Axline Computers manufactures personal computers at 2 plants. ...
Axline Computers manufactures personal computers at 2 plants. The Texas plant has 40 employees, the Hawaii plant has 20. A random sample of 10 employees is to be asked to fill out benefit questionnaires.
A. What is the probability that none of the employees in the sample work at the plant in Hawaii?
B. What is the probability that one of the employees in the sample works at the plant in Hawaii?
C. What is the probability that two or more in the sample work at the plant in Hawaii?
D. What is the probability that nine of the employees in the sample work at the plant in Hawaii?

Answer by stanbon(75887)

Axline Computers manufactures personal computers at 2 plants.
The Texas plant has 40 employees,
the Hawaii plant has 20.
---------------------------------------------
A random sample of 10 employees is to be asked to fill out benefit questionnaires.
A. What is the probability that none of the employees in the sample work at the plant in Hawaii?
---
Ans: 40C10/60C10 = 0.0112
--------------------------------------------
B. What is the probability that one of the employees in the sample works at the plant in Hawaii?
Ans: (20C1*40C9)/(60C10) = 0.0725
----------------------------------------------
C. What is the probability that two or more in the sample work at the plant in Hawaii?
Ans: 1 - [0.0112+0.0725] = 0.9163
----------------------------------------------
D. What is the probability that nine of the employees in the sample work at the plant in Hawaii?
Ans: [20C9*40C1]/60C10 = 0.00008911
===============
Cheers,
Stan H.